OpenStudy (lizz123):

Write an equation in slope intercept form the line passing through each pair of points. (4,6)(1,3) @jim_thompson5910

OpenStudy (lizz123):

\[m=\frac{ -3-6 }{1-4 }=\frac{ -9 }{ -3 }= 3\]

OpenStudy (lizz123):

After this I can choose any two points, correct?

jimthompson5910 (jim_thompson5910):

the -3 up top should be +3

jimthompson5910 (jim_thompson5910):

give it another shot and tell me what slope you get

jimthompson5910 (jim_thompson5910):

you should have \[\Large m = \frac{3-6}{1-4} = ??\]

OpenStudy (lizz123):

\[m=\frac{ 3 }{ -3 }\]

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

\[\Large m = \frac{3-6}{1-4} = \frac{-3}{-3} = 1\] you lost a sign when you computed 3-6

OpenStudy (lizz123):

oh yeah!!!

OpenStudy (lizz123):

Now, I am going to find b but i get confused doing this part

OpenStudy (lizz123):

I want to use the point (1,3)

OpenStudy (lizz123):

\[3=1(1)+b\]

jimthompson5910 (jim_thompson5910):

so far, so good

OpenStudy (lizz123):

\[3=1+b\]

jimthompson5910 (jim_thompson5910):

yep, also correct, now isolate b

OpenStudy (lizz123):